OpenStudy (anonymous):

Can anyone help me with the Sum, Difference and Product of two numbers equation? Suppose the functions of g and h are defined for all real numbers x as follows: g(x) = x -2 h(x) = 4x + 2 Write the expression for the following and evaluate: (g*h)(x) = ______ (g – h)(x) = ______ (g+h(1) = ______

OpenStudy (jdoe0001):

$$\bf g(x) = x -2\\ h(x) = 4x + 2\\ (g*h)(x) \implies g(x) \times h(x)\\ (g-h)(x) \implies g(x) - h(x)\\ (g+h)(1) \implies g(1) + h(1) $$ plug in your values and simplify :)

OpenStudy (anonymous):

(g*h)(x) = (x -2)*(4x+2) = ? (g - h)(x)= (x -2) - (4x+2) = ? (g+h(1)) = (x - 2) +(4x+2) then sub for 1 then solve

OpenStudy (anonymous):

Ok I am pretty sure this is incorrect: 1 hx -2 x-h -2 16

OpenStudy (anonymous):

what you get for (g*h)(x)?

OpenStudy (anonymous):

1 hx -2, I don't know this at all

OpenStudy (anonymous):

when you combine functions you just have to do what ever they tell you in this case you have to multiply them so lets take a look at this problem step-by-step. it asks us to multiply g*h so lets start bby writing the equation like this :\[(g*h)(x)=(x-2)*(4x+2)\] you could probb end here but i dont think you should since you can mulitply out the two poly. just use FOIL and you will have :\[(x-2)(4x+2)=4x^2+2x-8x-4= 4x^2 -6x -4\] which is your final answer if you have any question let me know i knida rushed this :(

OpenStudy (anonymous):

Ok, there is three sections to be answered, here is the equations: Suppose the functions of g and h are defined for all real numbers x as follows: g(x) = x -2 h(x) = 4x + 2 Write the expression for the following and evaluate: (g*h)(x) = ______ (g – h)(x) = ______ (g+h(1) = ______ What goes in each space after the equal sign?

OpenStudy (anonymous):

\[4x^2 - 6x - 4\]\[3x-4\]\[5\] in that order

OpenStudy (anonymous):

It looks like you are saying: The first answer is (g*h)(x) = (x-2)(4x+2) = 4x^2−6x−4 Right? Then we have (g-h)(x) = and (g+h(1) =

OpenStudy (anonymous):

How did you get 5 for (g+h) (1) ?

OpenStudy (anonymous):

i got 5 because its not asking us (g+h)(x) but its asking us (g+h)(1) so you have to replace every x with a 1 so here how i got it \[(g+h)(1) = (x-2) + (4x+2) = (4x + x)(-2 + 2) = 5x = 5(1) = 5\]

OpenStudy (anonymous):

hoped this help

OpenStudy (anonymous):

Looking at the book example it looks like, 4x^2 = 4*4=16*2=32 6x-4= 10 32*10=320

OpenStudy (anonymous):

I understand that you have written the expression, and I think you are right. It is then asking to evaluate it and the example shows multiplying the two expressions. I am not sure I did that correctly or not.

OpenStudy (anonymous):

well lets say this. if you have you (gh)(x) function lets take the first one in example 4x^2−6x−4 now that we know that if the problem ask us to evalute the problem f(2) just plug in the 2 to all the x's on that function

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

did i explain it well nuff for you to understand?

OpenStudy (anonymous):

I am working on it!

OpenStudy (anonymous):

kk :)

OpenStudy (anonymous):

Ok, I think this is it look at it and tell me what you think: (g*h)(x) = 64 (g – h)(x) = 4x^-6x-4 (g+h(1) = 4x^2+6x+4

OpenStudy (anonymous):

that should be 4x^2-6x-4

OpenStudy (anonymous):

how you get 64?

OpenStudy (anonymous):

By multiplying the x (2).

OpenStudy (anonymous):

ah i see alright :)